How many ways can you roll the number ten on 2 dice?
There are 36 permutations of 2 six-sided dice. Of those, 3 add up to ten; 4+6, 5+5, and 6+4. The probability of rolling ten is 3 in 36, or 1 in 18, or 0.05555...
What is the Probability of drawing two spades out of 52 cards?
The probability of drawing a spade in a standard 52 card deck is 13 in 52, or 1 in 4. The probability of drawing a second spade, assuming the first spade was not replaced back into the deck, is 12 in 51. The probability, then, of drawing two spades is the product of those two probabilities, or 12 in 204, or 1 in 17.
.9222
In the game of hearts the meaning of the queen of spades?
In the game of hearts, the meaning of the queen of spades is the equivalent of 13 hearts, or half the potential score. She is also known as "Dirty Gertie" because she can be played on someone, often at the last moment, with a low score, pushing their score (which is actually a negative score) up a significant amount, but without giving them enough to invert the score for that hand. She can also be left unplayed by the leader, who finds him/her self unable to get out of the lead, said player then getting the 13 points.
There are at least 11 green marbles in the bag.
The probability of selecting a 17 (or any number for that matter) is 1/20 or .05 or 5%.
This is always a fun type of question. The equation is summed up by n choose k, but it isn't easy to write that here....it looks something like this:
( n ) It is 26 choose 5 because you
( k )
( 26 )
( 5 )
That is, inputting the numbers: we need 5 of the 26 available red
cards. Now, how to solve?
( 26! )
( (26!-5!)! x 5! )
It equals:
65,780.
So, to conclude, there are a total of 65,780 ways of choosing 5 red cards from a 52 card deck.
Is the ace of spades tattoo racisit?
Yes it is. Some believe white women get the tattoo to show a preference in black men... What's it mean when a black man get the tattoo? Does he have a preference in black men?
It is the new symbol of the Black power movement.
White people (Especially Women) are just stupid.
The meaning of any tattoo is determined by the individual and their reason for getting it. What others think is their hang up and the most common symbolic meaning is simply the death card having no racial connotations whatsoever anyway.
124
Probability, P, of 70% or more correct (7 or more correct) is: P(7) + P(8) + P(9) + P(10). See the related link; N=10, P = 0.5, and K = 7, 8, 9, & 10. Therefore the probability is: .11719 + .04395 + .00977 + .00098 = .17189 or approximately 17.2% probability 7 or more correct.
7:1
What is the prbability of rolling a five on one die three times in a row?
The chance is one in 216 (6^-3).
The probability of rolling a five once is 1/6. Rolling a five again, on the same die or another, will still have a 1/6 chance. Therefore, the probability of the event occurring twice is 1/36 (1/6^2). Three times has a probability of 1/216 (1/6^3), and so on. It does not matter what die is used, as long as it has six sides.
The probability p of rolling a number x times consecutively on an s sided die is
p=s^-x
4/13
Let S be the sample space and E be the event of selecting 1 girl and 2 boys.
Then, n(S)= Number ways of selecting 3 students out of 25= 25C3 `=(25 x 24 x 23)(3 x 2 x 1)= 2300.
n(E)= (10C1 x 15C2)=10 x(15 x 14)(2 x 1)= 1050.
P(E) =n(E)=1050=21.n(S)230046
---------------------------------------------------------
O.K. but what is the answer to the question?
Two dice are rolled what is the probability of rolling at least one six?
First calculate the probability of not rolling a six - since there are 5 possibilities for each die, this is (5/6) x (5/6). Then calculate the complement (1 minus the probability calculated).
First calculate the probability of not rolling a six - since there are 5 possibilities for each die, this is (5/6) x (5/6). Then calculate the complement (1 minus the probability calculated).
First calculate the probability of not rolling a six - since there are 5 possibilities for each die, this is (5/6) x (5/6). Then calculate the complement (1 minus the probability calculated).
First calculate the probability of not rolling a six - since there are 5 possibilities for each die, this is (5/6) x (5/6). Then calculate the complement (1 minus the probability calculated).
Probability can be defined as a between the number of successes and number of possible outcomes?
ratio
The probability density function has what value in the interval between 20 and 28?
I think you left off some important information. Perhaps you can supply this information, to obtain assistance.
To calculate the probability or the chance of occurrence between two values, we calculate:
Pr{a < X < b} = F(b) - F(a)
where F(x) = cumulative probability distribution. The distribution requires certain known parameters. In the case of the Normal distribution, the mean and standard deviation are parameters.
In your particular case, a = 20 and b = 28.
If two fair dice were rolled, there would be 36 outcomes.
(1,1),(1,2),....,(6,6)
The maximum sum would be 12.
Therefore, the probability that the total number of spots shown is equal to 15 is zero.
The probability that both will be hopelessly romantic is .0081
.009^2 = .0081